HALF A PARABOLA

Mangho Ahuja

The locus of the midpoints of chords drawn from the vertex of the parabola y^{2} = 4px, is a parabola y^{2} = 2px.

Mangho Ahuja Department of Mathematics Southeast Missouri State University Cape Girardeau, MO 63701-4709 email: mahuja@semovm.semo.edu